# Unlocking the Potential of Classic GNNs for Graph-level Tasks: Simple Architectures Meet Excellence

### v1

@@ Abstract
Message-passing Graph Neural Networks (GNNs) are often criticized for their limited expressiveness, issues like over-smoothing and over-squashing, and challenges in capturing long-range dependencies, while Graph Transformers (GTs) are considered superior due to their global attention mechanisms. Literature frequently suggests that GTs outperform GNNs, particularly in graph-level tasks such as graph classification and regression. In this study, we explore the untapped potential of GNNs through an enhanced framework, GNN+, which integrates six widely used techniques: edge feature integration, normalization, dropout, residual connections, feed-forward networks, and positional encoding, to effectively tackle graph-level tasks. We conduct a systematic evaluation of three classic GNNs—GCN, GIN, and GatedGCN—enhanced by the GNN+ framework across 14 well-known graph-level datasets. Our results show that, contrary to the prevailing belief, classic GNNs excel in graph-level tasks, securing top three rankings across all datasets and achieving first place in eight, while also demonstrating greater efficiency than GTs. This highlights the potential of simple GNN architectures, challenging the belief that complex mechanisms in GTs are essential for superior graph-level performance. Our source code is available at https://github.com/LUOyk1999/tunedGNN-G.
@@ 1 Introduction
Graph machine learning addresses both graph-level tasks and node-level tasks, as illustrated in Figure 1. These tasks fundamentally differ in their choice of the basic unit for dataset composition, splitting, and training, with graph-level tasks focusing on the entire graph, while node-level tasks focus on individual nodes. Graph-level tasks (Dwivedi et al., 2023; Hu et al., 2020; Luo et al., 2023b, a) often involve the classification of relatively small molecular graphs in chemistry (Morris et al., 2020) or the prediction of protein properties in biology (Dwivedi et al., 2022). In contrast, node-level tasks typically involve large social networks (Tang et al., 2009) or citation networks (Yang et al., 2016), where the primary goal is node classification. This distinction in the fundamental unit of dataset leads to differences in methodologies, training strategies, and application domains.
@@
Figure 1: Differences between graph-level and node-level tasks.
A prevalent perspective is that Graph Transformers (GTs) (Müller et al., 2023; Min et al., 2022; Hoang et al., 2024), as an alternative to GNNs, leverage global attention mechanisms that enable each node to attend to all others (Yun et al., 2019; Dwivedi & Bresson, 2020), effectively modeling long-range interactions and addressing issues such as over-smoothing, over-squashing, and limited expressiveness (Kreuzer et al., 2021; Ying et al., 2021; Zhang et al., 2023; Luo et al., 2023c, 2024b). However, the quadratic complexity of global attention mechanisms limits the scalability of GTs in large-scale, real-world applications (Behrouz & Hashemi, 2024; Sancak et al., 2024; Ding et al., 2024). Moreover, it has been noted that many state-of-the-art GTs (Chen et al., 2022; Rampášek et al., 2022; Shirzad et al., 2023; Ma et al., 2023) still rely—either explicitly or implicitly—on the message passing mechanism of GNNs to learn local node representations, thereby enhancing performance.
Recent studies (Luo et al., 2024a, 2025a, 2025b) have shown that, contrary to common belief, classic GNNs such as GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2018), and GraphSAGE (Hamilton et al., 2017) can achieve performance comparable to, or even exceeding, that of state-of-the-art GTs for node-level tasks. However, a similar conclusion has not yet been established for graph-level tasks. While Tönshoff et al. (2023) conducted pioneering research demonstrating that tuning a few hyperparameters can significantly enhance the performance of classic GNNs, their results indicate that these models still do not match the overall performance of GTs. Furthermore, their investigation is limited to the Long-Range Graph Benchmark (LRGB) (Dwivedi et al., 2022). This raises an important question: “Can classic GNNs also excel in graph-level tasks?”
@@
We systematically evaluate 3 classic GNNs—GCN (Kipf & Welling, 2017), GIN (Xu et al., 2018), and GatedGCN (Bresson & Laurent, 2017)—enhanced by the GNN+ framework across 14 well-known graph-level datasets from GNN Benchmark (Dwivedi et al., 2023), LRGB (Dwivedi et al., 2022), and OGB (Hu et al., 2020). The results demonstrate that the enhanced versions of classic GNNs match or even outperform state-of-the-art (SOTA) GTs, achieving rankings in the top three, including first place in eight datasets, while exhibiting superior efficiency. These findings provide a positive answer to the previously posed question, suggesting that the true potential of GNNs for graph-level applications has been previously underestimated, and the GNN+ framework effectively unlocks this potential while addressing their inherent limitations. Our ablation study also highlights the importance of each technique used in GNN+ and offers valuable insights for future research.
@@ 4 Assessment: Experimental Setup
Baselines. Our main focus lies on classic GNNs: GCN (Kipf & Welling, 2017), GIN (Xu et al., 2018; Hu et al., 2019), GatedGCN (Bresson & Laurent, 2017), the SOTA GTs: GT (2020), GraphTrans (2021), SAN (2021), Graphormer (2021), SAT (2022), EGT (2022), GraphGPS (2022; 2023), GRPE (2022), Graphormer-URPE (2022), Graphormer-GD (2023), Specformer (2023), LGI-GT (2023), GPTrans-Nano (2023b), Graph ViT/MLP-Mixer (2023), NAGphormer (2023a), DIFFormer (2023), MGT (2023), DRew (2023), Exphormer (2023), GRIT (2023), GRED (2024), GEAET (2024), Subgraphormer (2024), TIGT (2024), GECO (2024), GPNN (2024), Cluster-GT (2024a), and the SOTA graph state space models (GSSMs): GMN (2024), Graph-Mamba (2024), GSSC (2024b). Furthermore, various other GTs exist in related surveys (Hoang et al., 2024; Shehzad et al., 2024; Müller et al., 2023), empirically shown to be inferior to the GTs we compared against for graph-level tasks. We report the performance results of baselines primarily from (Rampášek et al., 2022; Tönshoff et al., 2023), with the remaining obtained from their respective original papers or official leaderboards whenever possible, as those results are obtained by well-tuned models.
@@ 5.2 Ablation Study
For large-scale datasets such as CIFAR 10, COCO-SP, and the OGB datasets, removing normalization leads to significant performance drops. Specifically, on ogbg-ppa, which has 158,100 graphs, ablating normalization results in an accuracy drop of around 15% for three classic GNNs. This result is consistent with Luo et al. (2024a), who found that normalization is more important for GNNs in node classification on large graphs. In such datasets, where node feature distributions are more complex, normalizing node embeddings is essential for stabilizing the training process.
@@
Our analysis highlights the crucial role of dropout in maintaining the performance of classic GNNs on GNN Benchmark and LRGB and large-scale OGB datasets, with its ablation causing significant declines—for instance, an 8.8% relative decrease for GatedGCN+ on CIFAR-10 and a 20.4% relative decrease on PascalVOC-SP. This trend continues in large-scale OGB datasets, where removing dropout results in a 5–13% performance drop across 3 classic GNNs on ogbg-molpcba. Notably, 97% of the optimal dropout rates are $\leq$ 0.2, and 64% are $\leq$ 0.1, indicating that a very low dropout rate is both sufficient and optimal for graph-level tasks. Interestingly, this finding for graph-level tasks contrasts with Luo et al. (2024a)’s observations for node-level tasks, where a higher dropout rate is typically required.
@@ 6 Conclusion
This study highlights the often-overlooked potential of classic GNNs in tacking graph-level tasks. By integrating six widely used techniques into a unified GNN+ framework, we enhance three classic GNNs for graph-level tasks. Evaluations on 14 benchmark datasets reveal that, these enhanced GNNs match or outperform GTs, while also demonstrating greater efficiency. These findings challenge the prevailing belief that GTs are inherently superior, reaffirming the capability of simple GNN structures as powerful models.
Impact Statements
This paper presents work whose goal is to advance the field of Graph Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.

### v2

# Can Classic GNNs Be Strong Baselines for Graph-level Tasks? Simple Architectures Meet Excellence

@@ Abstract
[…]
@@
Figure 1: Differences between graph-level and node-level tasks.
@@ 1 Introduction
Graph machine learning addresses both graph-level tasks and node-level tasks, as illustrated in Figure 1. These tasks fundamentally differ in their choice of the basic unit for dataset composition, splitting, and training, with graph-level tasks focusing on the entire graph, while node-level tasks focus on individual nodes. Graph-level tasks (Dwivedi et al., 2023; Hu et al., 2020; Luo et al., 2023b, a) often involve the classification of relatively small molecular graphs in chemistry (Duvenaud et al., 2015; Morris et al., 2020) or the prediction of protein properties in biology (Dwivedi et al., 2022). In contrast, node-level tasks typically involve large social networks (Tang et al., 2009) or citation networks (Yang et al., 2016), where the primary goal is node classification. This distinction in the fundamental unit of dataset leads to differences in methodologies, training strategies, and application domains.
@@
A prevalent perspective is that Graph Transformers (GTs) (Müller et al., 2023; Min et al., 2022; Luo et al., 2024c, a), as an alternative to GNNs, leverage global attention mechanisms that enable each node to attend to all others (Yun et al., 2019; Dwivedi & Bresson, 2020), effectively modeling long-range interactions and addressing issues such as over-smoothing, over-squashing, and limited expressiveness (Rampášek et al., 2022; Zhang et al., 2023). Consequently, GTs have become the preferred method for graph-level tasks, especially in the context of small molecular graphs (Ma et al., 2023). However, the quadratic complexity of global attention mechanisms limits the scalability of GTs in large-scale, real-world applications (Behrouz & Hashemi, 2024; Sancak et al., 2024; Ding et al., 2024). Moreover, it has been noted that many state-of-the-art GTs (Chen et al., 2022; Rampášek et al., 2022; Shirzad et al., 2023; Ma et al., 2023) still rely—either explicitly or implicitly—on the message passing mechanism of GNNs to learn local node representations, thereby enhancing performance.
Recent studies (Luo et al., 2024b, 2025a, 2025b) have demonstrated that classic GNNs such as GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2018), and GraphSAGE (Hamilton et al., 2017) can achieve performance comparable to, or even exceeding, that of state-of-the-art GTs for node-level tasks. However, a similar conclusion has not yet been established for graph-level tasks. While Tönshoff et al. (2023) conducted pioneering research demonstrating that tuning a few hyperparameters can significantly enhance the performance of classic GNNs, their results indicate that these models still do not match the overall performance of GTs. Furthermore, their investigation is limited to the Long-Range Graph Benchmark (LRGB) (Dwivedi et al., 2022). This raises an important question: “Can classic GNNs also excel in graph-level tasks?”
@@
We systematically re-evaluate 3 classic GNNs—GCN (Kipf & Welling, 2017), GIN (Xu et al., 2018), and GatedGCN (Bresson & Laurent, 2017)—enhanced by the GNN+ framework across 14 well-known graph-level datasets from GNN Benchmark (Dwivedi et al., 2023), LRGB (Dwivedi et al., 2022), and OGB (Hu et al., 2020). The results show that the enhanced versions of classic GNNs match or even outperform state-of-the-art (SOTA) GTs, consistently securing top-three rankings and achieving first place in eight datasets. Moreover, they exhibit superior efficiency, running several times faster than GTs on many datasets. These findings provide a positive answer to the previously posed question, suggesting that the true potential of GNNs for graph-level applications has been underestimated. The GNN+ framework effectively unlocks this potential while addressing inherent limitations.
Furthermore, in our ablation study, we highlight the importance of each technique integrated into GNN+. Specifically,
  1. (1)
    Edge features are particularly beneficial in molecular and image superpixel datasets, where they encode crucial domain-specific information.
  2. (2)
    Normalization becomes more crucial as the scale of datasets increases.
  3. (3)
    Dropout benefits most graph-level datasets, where a modest dropout rate proves both sufficient and optimal.
  4. (4)
    Residual connections are consistently essential, except in shallow GNNs applied to small graphs.
  5. (5)
    FFN is especially pivotal for simpler models, such as GCN, in graph-level tasks.
  6. (6)
    Positional encoding plays a more critical role in small-scale datasets compared to large-scale ones.
These findings underscore the nuanced roles each technique plays in enhancing the performance of classic GNNs, guiding future explorations in GNN design and application.
@@ 4 Assessment: Experimental Setup
Baselines. Our main focus lies on classic GNNs: GCN (Kipf & Welling, 2017), GIN (Xu et al., 2018; Hu et al., 2019), GatedGCN (Bresson & Laurent, 2017), the SOTA GTs: GT (2020), GraphTrans (2021), SAN (2021), Graphormer (2021), SAT (2022), EGT (2022), GraphGPS (2022; 2023), GRPE (2022), Graphormer-URPE (2022), Graphormer-GD (2023), Specformer (2023), LGI-GT (2023), GPTrans-Nano (2023b), Graph ViT/MLP-Mixer (2023), NAGphormer (2023a), DIFFormer (2023), MGT (2023), DRew (2023), Exphormer (2023), GRIT (2023), GRED (2024), GEAET (2024), Subgraphormer (2024), TIGT (2024), GECO (2024), GPNN (2024), Cluster-GT (2024a), and the SOTA graph state space models (GSSMs): GMN (2024), Graph-Mamba (2024), GSSC (2024b). Furthermore, various other GTs exist in related surveys (Hoang et al., 2024; Shehzad et al., 2024; Müller et al., 2023), empirically shown to be inferior to the GTs we compared against for graph-level tasks. We report the performance results of baselines primarily from (Rampášek et al., 2022; Tönshoff et al., 2023; Ma et al., 2023; Wang et al., 2024), with the remaining obtained from their respective original papers or official leaderboards whenever possible, as those results are obtained by well-tuned models.
@@ 5.2 Ablation Study
For large-scale datasets such as CIFAR 10, COCO-SP, and the OGB datasets, removing normalization leads to significant performance drops. Specifically, on ogbg-ppa, which has 158,100 graphs, ablating normalization results in an accuracy drop of around 15% for three classic GNNs. This result is consistent with Luo et al. (2024b), who found that normalization is more important for GNNs in node classification on large graphs. In such datasets, where node feature distributions are more complex, normalizing node embeddings is essential for stabilizing the training process.
@@
Our analysis highlights the crucial role of dropout in maintaining the performance of classic GNNs on GNN Benchmark and LRGB and large-scale OGB datasets, with its ablation causing significant declines—for instance, an 8.8% relative decrease for GatedGCN+ on CIFAR-10 and a 20.4% relative decrease on PascalVOC-SP. This trend continues in large-scale OGB datasets, where removing dropout results in a 5–13% performance drop across 3 classic GNNs on ogbg-molpcba. Notably, 97% of the optimal dropout rates are $\leq$ 0.2, and 64% are $\leq$ 0.1, indicating that a very low dropout rate is both sufficient and optimal for graph-level tasks. Interestingly, this finding for graph-level tasks contrasts with Luo et al. (2024b)’s observations for node-level tasks, where a higher dropout rate is typically required.
@@ 6 Conclusion
This study highlights the often-overlooked potential of classic GNNs in tacking graph-level tasks. By integrating six widely used techniques into a unified GNN+ framework, we enhance three classic GNNs (GCN, GIN, and GatedGCN) for graph-level tasks. Evaluations on 14 benchmark datasets reveal that, these classic GNNs match or outperform SOTA GTs, while also demonstrating greater efficiency. These findings challenge the prevailing belief that GTs are inherently superior, reaffirming the capability of simple GNN structures as powerful models.
Acknowledgments
We sincerely thank the anonymous area chair for their thoughtful evaluation and strong support, which gave us the opportunity to share this work with the community. We are also grateful to all anonymous reviewers for their constructive and insightful feedback. We sincerely thank Yiwen Sun for her invaluable help with refining the writing and to Lu Fan for her help with creating Figure 1.
This work received support from National Key R&D Program of China (2021YFB3500700), NSFC Grant 62172026, National Social Science Fund of China 22&ZD153, the Fundamental Research Funds for the Central Universities, State Key Laboratory of Complex & Critical Software Environment (SKLCCSE), and the HK PolyU Grant P0051029. Lei Shi is with School of Computer Science and Engineering, Beihang University, and the State Key Laboratory of Complex & Critical Software Environment.
Impact Statement
This paper presents an empirical benchmarking study that re-evaluates the performance of classic GNNs on graph-level tasks. By systematically integrating widely used techniques into a unified framework (GNN+), the study demonstrates that classic GNNs can achieve performance comparable to SOTA models. The findings challenge common assumptions in the field and provide practical guidance for future model selection and design. As this is a benchmarking effort, we do not foresee any immediate societal consequences that must be specifically highlighted.
@@ A.3 Hyperparameters and Reproducibility
Appendix B Additional Experimental Results
B.1 Hyperparameter Sensitivity Analysis
To further understand the robustness of GNN+, we conduct a comprehensive sensitivity analysis focusing on two key hyperparameters: the number of layers and the dropout rate.
Layer Depth.
We investigate the impact of the number of network layers on three datasets: PATTERN and CLUSTER, and PascalVOC-SP. As shown in Figure 3, GNN+ achieves strong and stable performance across a wide depth range. On PATTERN, both GCN+ and GatedGCN+ reach peak performance at 12 layers. In contrast, on CLUSTER and PascalVOC-SP, performance continues to improve as the number of layers increases, suggesting that GNN+ is capable of leveraging deeper architectures when appropriate.
Importantly, when residual connections are disabled (bottom row), we observe significant performance degradation as the number of layers increases, especially beyond 8 layers. This is most pronounced on CLUSTER and PascalVOC-SP, where over-smoothing and over-squashing effects become severe in the absence of residual pathways. These results confirm that residual connections play a crucial role in stabilizing deep message propagation in GNN+.
Figure 3: Experiments on the impact of the number of layers in GNN+ on PATTERN, CLUSTER, and PascalVOC-SP. Top row: performance with residual connections; bottom row: performance without residual connections.
Dropout Rate.
We conduct additional experiments to analyze the effect of dropout rates (ranging from 0.0 to 0.5) on four datasets: CLUSTER, PascalVOC-SP, Peptides-struct, and Peptides-func. As shown in Figure 4, the optimal dropout rate lies in the range of 0.1 to 0.2. Performance degrades significantly when the dropout rate reaches 0.3 or higher across all datasets. For instance, on Peptides-struct, the MAE increases sharply beyond 0.2 dropout, and on CLUSTER, accuracy collapses at 0.5. This suggests that excessive dropout disrupts the message-passing process in graph-level tasks.
Figure 4: Sensitivity analysis of dropout rates in GNN+ on CLUSTER, PascalVOC-SP, Peptides-struct, and Peptides-func.
Binary Hyperparameters.
For components that can be enabled or disabled, such as edge features, normalization, residual connections, FFN, and positional encoding, we refer readers to the ablation results in Tables 5 and 6.
B.2 Performance of a Fixed GNN+ Configuration
We conduct an additional experiment using a fixed architecture of GNN+, in which all six techniques are integrated. This model referred to as GNN+ (fixed) is applied uniformly across datasets without dataset-specific tuning. The results are summarized in Table 14. The performance of GNN+ (fixed) is comparable to the best configurations reported in the main paper, further demonstrating the model’s practicality and robustness without requiring extensive tuning.
Table 14: Performance of GNN+ with fixed configuration (i.e., all six techniques integrated) across 9 datasets.
PascalVOC-SP COCO-SP MalNet-Tiny ogbg-molpcba ogbg-code2 MNIST CIFAR10 PATTERN CLUSTER
GCN+ 0.3357 $\pm$ 0.0087 0.2733 $\pm$ 0.0041 0.9354 $\pm$ 0.0045 0.2721 $\pm$ 0.0046 0.1787 $\pm$ 0.0026 98.382 $\pm$ 0.095 69.824 $\pm$ 0.413 87.021 $\pm$ 0.095 77.109 $\pm$ 0.872
GCN+ (fixed) 0.3341 $\pm$ 0.0055 0.2716 $\pm$ 0.0034 0.9235 $\pm$ 0.0060 0.2694 $\pm$ 0.0059 0.1784 $\pm$ 0.0029 98.257 $\pm$ 0.063 69.436 $\pm$ 0.265 87.021 $\pm$ 0.095 76.352 $\pm$ 0.757
GatedGCN+ 0.4263 $\pm$ 0.0057 0.3802 $\pm$ 0.0015 0.9460 $\pm$ 0.0057 0.2981 $\pm$ 0.0024 0.1896 $\pm$ 0.0024 98.712 $\pm$ 0.137 77.218 $\pm$ 0.381 87.029 $\pm$ 0.037 79.128 $\pm$ 0.235
GatedGCN+ (fixed) 0.4204 $\pm$ 0.0061 0.3774 $\pm$ 0.0028 0.9450 $\pm$ 0.0045 0.2981 $\pm$ 0.0024 0.1889 $\pm$ 0.0018 98.712 $\pm$ 0.137 77.218 $\pm$ 0.381 87.029 $\pm$ 0.037 79.128 $\pm$ 0.235
B.3 Additional Baselines
Comparison with Fast Linear Models.
To assess the performance-efficiency trade-off, we compare GCN+ with two recent linear-time graph classification baselines, LDP (Cai & Wang, 2018) and HRN (Wu et al., 2022), on both small-scale and large-scale datasets, using the same training protocols and hyperparameter search spaces.
We first evaluate all models on five small-scale graph classification datasets used in (Cai & Wang, 2018; Wu et al., 2022): IMDB-B, IMDB-M, COLLAB, MUTAG, and PTC. These datasets contain between 300 and 5,000 graphs. As shown in Table 15, GCN+ achieves accuracy comparable to or exceeding that of LDP and HRN on all datasets. Given the small scale, training times are negligible across models.
Table 15: Accuracy (%) on small-scale graph classification datasets.
IMDB-B IMDB-M COLLAB MUTAG PTC
# graphs 1000 1500 5000 188 344
LDP 75.4 50.0 78.1 90.3 64.5
HRN 77.5 52.8 81.8 90.4 65.7
GCN+ 76.9 52.3 80.9 90.1 66.6
To examine scalability, we evaluate all models on two OGB datasets: ogbg-molhiv and ogbg-molpcba. While LDP and HRN exhibit lower training costs, their predictive performance is significantly lower than GCN+. Table 16 summarizes this trade-off, demonstrating the importance of benchmarking on large datasets for assessing practical utility.
Table 16: Performance and training cost on OGB molecular property prediction datasets.
ogbg-molhiv ogbg-molpcba
AUROC $\uparrow$ Time (s/epoch) $\downarrow$ Avg. Precision $\uparrow$ Time (s/epoch) $\downarrow$
LDP 0.7121 $\pm$ 0.0105 5 0.1243 $\pm$ 0.0031 31
HRN 0.7587 $\pm$ 0.0147 9 0.2274 $\pm$ 0.0043 68
GCN+ 0.8012 $\pm$ 0.0124 16 0.2721 $\pm$ 0.0046 91
Comparison with SOTA GNNs for Over-smoothing and Over-squashing.
We further compare GCN+ with two recently proposed models designed to mitigate over-smoothing and over-squashing: Multi-Track GCN (MTGCN) (Pei et al., 2024) and Cooperative GNN (CO-GNN) (Finkelshtein et al., 2024). All models are evaluated on Peptides-func and Peptides-struct using consistent training protocols and hyperparameter search spaces.
For CO-GNN, we tune the message-passing scheme (SUMGNN, MEANGNN, GCN, GIN, GAT) as recommended by the original paper. For MTGCN, we optimize the number of message-passing stages (from 1 to 4). Results are shown in Table 17. GCN+ consistently outperforms both models on both datasets, confirming its effectiveness in addressing these limitations while retaining architectural simplicity.
Table 17: Comparison with models addressing over-smoothing and over-squashing.
Peptides-func Peptides-struct
MTGCN 0.6936 $\pm$ 0.0089 0.2461 $\pm$ 0.0019
CO-GNN 0.7012 $\pm$ 0.0106 0.2503 $\pm$ 0.0025
GCN+ 0.7261 $\pm$ 0.0067 0.2421 $\pm$ 0.0016
B.4 Visualization
To gain further insights into the representational quality of GNN+, we visualize the learned graph embeddings using t-SNE. Specifically, we compare embeddings generated by GCN and GCN+ on Peptides-func.
Figure 5 shows the resulting t-SNE projections. We observe that the embeddings produced by GCN+ exhibit more distinct class separation and greater inter-class margins compared to those produced by the vanilla GCN.
Figure 5: t-SNE visualization of graph-level embeddings learned by GCN (left) and GCN+ (right). Colors denote different classes.
Figure 6: Visualization of GraphGPS attention scores across all nodes within the same molecular graph. The left side illustrates the molecular structure. The center column displays the attention scores of a node (highlighted with green borders) learned by the classic GraphGPS, while the right column showcases the 3-hop neighborhood of the node. Notably, the majority of nodes attend predominantly to two structural motifs: the five-membered ring (pyrazoline, top right) and the benzene ring (bottom). In contrast, functionally significant substructures—such as the C–O–N=N group (top left) and the nitro group (–NO2, bottom left)—receive minimal attention.
B.5 Static Attention Problem of GTs
To gain insight into why GNNs can outperform GTs in graph-level tasks, we analyze the attention patterns in several representative GT architectures. We identified a potential limitation, which we call the static attention problem: the attention scores appear to be globally consistent across all nodes in the graph, irrespective of the query node.
As a concrete example, we analyze GraphGPS on a misclassified molecule from the ZINC dataset (Figure 6). As illustrated, nearly all nodes primarily focus on two structures: the five-membered pyrazoline ring in the top right and the benzene ring at the bottom. In contrast, functionally significant substructures, such as the C–O–N=N group in the top left and the nitro group (–NO2) in the bottom left, receive minimal attention. This query-invariant attention pattern results in insufficient sensitivity to subgraph structures, which adversely affects prediction accuracy.
In contrast, message-passing GNNs perform node-specific aggregation, allowing the model to capture diverse local substructures more effectively. This is beneficial for graph prediction tasks, where node representations are aggregated (e.g., via global pooling) into a global graph embedding. When nodes encode meaningful subgraph patterns, the resulting graph representation becomes more informative and discriminative.
Appendix C Further Discussion
C.1 Clarification on GNN+ and Classic GNNs
GNN+ enhances classic message-passing GNNs by integrating widely adopted architectural techniques including edge features, normalization, dropout, residual connections, FFN, and positional encoding. These components are not novel or specific to any one architecture; rather, they have been widely used across the literature on classic GNNs:
  • Residual connections, dropout, and normalization were included in the original GCN (Kipf & Welling, 2017).
  • Edge feature is already incorporated in the early general MPNN framework (Gilmer et al., 2017), which forms the foundation of many classic GNNs such as GatedGCN (Bresson & Laurent, 2017).
  • The use of an MLP/FFN after message passing was introduced in the GIN paper (Xu et al., 2018).
  • Positional encoding has been employed in prior GNN works  (Dwivedi et al., 2021).
GNN+ unifies these standard components within a consistent framework while preserving the core message-passing mechanism. Our results show that classic GNNs, when equipped with such enhancements, can achieve performance comparable to SOTA models on graph-level tasks. We view GNN+ as a practical and faithful extension that unlocks the potential of classic GNNs. In this sense, GNN+ is not a departure from the message-passing paradigm, but rather a natural evolution that reinforces its core strengths.
C.2 Related Works
Our work relates closely to recent efforts in benchmarking GNNs. As node-level benchmarks have received significantly more attention, several studies such as (Shchur et al., 2018; Wang et al., 2021; Lv et al., 2021) have examined the reliability of existing evaluation protocols, revealing that many newly proposed architectures fail to outperform simple baselines under fair conditions. Building on this line of inquiry, Luo et al. (2024b) demonstrated that classic GNNs can match or even surpass SOTA models across 18 widely used node classification datasets.
In contrast, our study targets graph-level tasks. Errica et al. (2020) highlights the importance of rigorous evaluation protocols and shows that many claimed improvements vanish under fair settings. Hu et al. (2020) introduces the OGB benchmark to standardize large-scale graph evaluations. Dwivedi et al. (2023) introduces a diverse suite of graphs for fair architectural comparisons under fixed parameter budgets, and Dwivedi et al. (2022) proposes the LRGB to assess models on long-range dependency tasks. Tönshoff et al. (2023) further investigates LRGB and shows that classic GNNs can close the gap with GTs under proper tuning. Grötschla et al. (2024) offers a systematic study of positional encoding. However, these studies have been constrained in both scope and comprehensiveness, primarily due to the limited number and diversity of datasets used, as well as an incomplete examination of GNN hyperparameters.
C.3 Limitations and Future Work
While our study presents strong empirical evidence that enhanced classic GNNs can serve as strong baselines for graph-level tasks, it remains primarily an empirical benchmarking effort. As such, we do not provide a theoretical analysis of why these classic GNNs can outperform GTs. This omission represents a key limitation of the current work.
A deeper theoretical understanding, for example in terms of expressive power, the mechanisms underlying over-smoothing and over-squashing, and the interplay between architectural components, could offer valuable insights into the design of more effective GNN variants. In future work, we aim to formally characterize the conditions under which GNN+ achieves superior performance and to develop principled analyses of its expressiveness and representational behavior.